C = 450 + 0.4Y
I = 350

G = 150
X = 70
Z = 35 + 0.1Y
T = 0.15Y
Yf = 1550




)

Q.2.3 Calculate the equilibrium level of income
(Hint: use the multiplier
1
Expert's answer
2020-12-06T18:25:10-0500

Y = C + I + G + X – Z

Y = 450 + 0.4Y + 350 + 150 + 70 – 35 – 0.1Y

Y = 985 + 0.4(Y – T) – 0.1Y

Y = 985 + 0.4(Y – 0.15Y) – 0.1Y

Y = 985 + 0.24Y

Y – 0.24Y = 985

0.76Y = 985

Equilibrium level of income = 1296